If 15.0 mL of phosphoric acid completely neutralizes 38.5 mL of 0.150 mol/L calcium hydroxide, what is the concentration of the phosphoric acid?


1
Expert's answer
2021-06-11T01:58:39-0400

2H3PO4 + 3Ca(OH)2 = Ca3PO4 + 6H2O

According to the equation, n (H3PO4) = 2/3 n (Ca(OH)2).

CM = n / V

n (Ca(OH)2) = 0.150 x 0.0385 = 0.006 mol

n (H3PO4) = 2/3 x 0.006 = 0.004 mol

CM (H3PO4) = 0.004 / 0.015 = 0.26 M